# The dark side of the forces: assessing non-conservative force models for atomistic machine learning

### 19 Feb 2025 v2

@@ 4.5 Equilibrium properties in the NVT ensemble
Judging by the average temperature $\langle T\rangle$ (Table 2), it is relatively easy to control the non-conservative behavior using a white-noise (WN) Langevin thermostat. However, strong couplings $\tau$ (the time scale over which the thermostat interferes with atomic motion) are needed, as even at 100 fs the average temperature is significantly off the target value. We discuss how these deviations in the equilibrium temperature affect structural properties of water in Section F.3. The upshot is that for accurate models and strong thermostatting, the effects are small but noticeable. Furthermore, the strong Langevin thermostatting reduces the sampling efficiency, and so even with a respectable 1 ns trajectory, many simple structural averages are not fully converged.
@@ 4.6 Sampling efficiency and time-dependent properties
Aggressive Langevin dynamics is bound to dramatically change time-dependent properties, and in particular to reduce the diffusion coefficient – so that longer trajectories are needed to collect statistically independent atomic configurations. This slow-down is apparent when looking at the velocity correlation spectra (Figure 3). In the weak coupling regime (WN, $\tau=1000$ fs) there is a (small) increase in diffusion coefficient relative to the reference, because of the unphysically higher temperature, while the high-frequency peaks corresponding to stretching and bending are only weakly perturbed. Stronger couplings alter the dynamics dramatically, and reduce the diffusion coefficient (and hence the efficiency in sampling slow, collective motion) by a factor of about 1.5 ($\tau=100$ fs) and 5 ($\tau=10$ fs), negating the inference speed-up for a non-conservative model – while making it impossible to accurately evaluate any time-dependent property.
Figure 3: Velocity power spectrum $\hat{c}_{vv}(\omega)$, for different PET models and thermostat types. See Table 2 for an explanation of the acronyms used in the legend.
@@ 4.7 Geometry optimization
MD performed at very low temperature can be regarded as a form of geometry optimization. Our observations from MD suggest that sufficiently accurate non-conservative models, should be able to reach reasonable, low-energy structures. We restrict ourselves to optimization algorithms based only on gradients, to avoid the complication of using inconsistent energies, as discussed in Section 3.3, comparing the FIRE (Bitzek et al., 2006) algorithm – that is similar in spirit to zero-temperature MD – and LBFGS (Liu & Nocedal, 1989) – a quasi-Newton algorithm that uses an approximation of the Hessian to accelerate convergence. Comparing different models on the task of optimizing a water snapshot from a MD configuration (Figure 6) shows that inaccurate non-conservative models, such as SOAP-BPNN-NC, fail catastrophically at geometry optimization, while more accurate models, such as PET-NC and ORB, can reach a locally stable configuration, especially using FIRE. We note, however, that non-conservative models are less stable when used with a Hessian-based method, with large fluctuations in the residual force that make it hard to define a stopping criterion. On a practical level, non-conservative forces are bound to make geometry optimization more fragile, and to require careful choice of the minimization algorithm and its convergence parameters, as we observe in Appendix E when comparing different general-purpose models.
@@ 4.8 Non-conservative forces as accelerators
While we have observed that conservative MLIPs are better suited for practical simulations, we suggest that hybrid models, which additionally support direct, non-conservative, force predictions, can be used for faster inference and training. Such models can be obtained by training both force heads jointly, as demonstrated in the PET-M model, or, more efficiently, by first training a non-conservative model and then fine-tuning its energy head to yield conservative forces. As shown in Figure 4 and further discussed in Appendix G, this leads to the accuracy and physical correctness of conservative models at highly reduced training time.
Figure 4: Training curves for PET-C and PET-NC models, and the fine-tuning of a hybrid PET-M model initialized (epoch marked with arrow) from the potential energy head of the PET-NC model.
In simulations, one can then use the conservative forces of such a hybrid model for validation, error monitoring and correction, and the direct forces for faster inference. A good example is to use multiple time stepping techniques for molecular dynamics, where the non-conservative forces are used to integrate the equations of motion, and the conservative forces are applied every few steps as a correction. This is further explained in Appendix H; results using this technique are shown in Table 2, Figure 2 and in Appendix H and demonstrate that 1/8th of conservative force evaluations are sufficient to obtain results that are essentially indistinguishable from fully conservative ones.
@@ 5 Discussion
Code and data required to reproduce the results in this work are available on Zenodo at https://zenodo.org/records/14778891. More details on the software used are available in Appendix I.
@@
Acknowledgements
The authors would like to thank Federico Grasselli for stimulating discussion, and Rafael Gomez-Bombarelli for a coffee-break conversation which inspired us to look into this problem. ML and MC acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme Grant No. 101001890-FIAMMA. FB and MC acknowledge support from the NCCR MARVEL, funded by the Swiss National Science Foundation (SNSF, grant number 182892) and from the Swiss Platform for Advanced Scientific Computing (PASC).
@@ Appendix B Theoretical computational cost
Appendix C Models and architectures
@@
ORB Non-equivariant, non-conservative model, trained on the Alexandria and MPtrj datasets.
@@
C.1 Timings of general-purpose models
@@
C.2 Accuracy and timings of water models
@@
Appendix D Non-conservative work
@@
Appendix E Geometry optimization
E.1 Optimization trajectories
@@
We discuss first the results for the FIRE algorithm (left panel in Figure 6). Despite the very different level of accuracy, the conservative SOAP-BPNN model and PET converge to a similar configuration (with PET forces saturating at about $10^{-3}$ eV/Å, due to numerical precision). The non-conservative PET-NC model also converges to a similar structure, indicating that a sufficiently accurate non-conservative model can indeed be used with a gradient-based structural optimization algorithm. The non-conservative SOAP-BPNN model, however, displays a catastrophic mode of failure, with the force never decreasing below 0.1 eV/Å, and the trajectory drifting off in a different direction without never reaching a stable state. ORB converges to a very different structure than the other models, which might be due to the different reference DFT functional used for training.
@@ E.2 Failure rates
Appendix F Additional molecular dynamics results
F.1 NVE temperature profiles
@@
We also perform NVE MD trajectories on a wider range of chemical systems using the general-purpose potentials MACE-MP, SevenNet, Equiformer and ORB. Fig. 8 shows that the non-conservative models ORB and equiformer exhibit a strong kinetic energy gain during the simulation of graphene, amorphous carbon and FCC aluminium, with ORB exhibiting an explosive behavior for graphene (possibly because it is outside the training set). The conservative models exhibit stable temperature fluctuations, even for amorphous carbon where the high distortion of the structure leads to higher initial energy and broader kinetic energy fluctuations.
F.2 NVT
@@
Table 7: Average kinetic temperatures for different models and thermostatting schemes. For SVR, we show also the separate kinetic temperature computed only for hydrogen and oxygen atoms.
@@
F.3 Structural correlations
@@
Appendix G Conservative fine-tuning
@@
Appendix H Multiple time stepping
@@
where $\mathbf{q}$ and $\mathbf{p}$ indicate the vectors containing positions and momenta of all particles, and $\mathbf{m}$ the masses. Forces have to be recomputed if they are needed after a position update. In the limit of ideal time scale separation, the resulting dynamics (and the sampled ensemble for constant-temperature simulations) is consistent with $V_{\text{slow}}$, at a much-reduced cost. See also Morrone et al. (2011) for a discussion of the failure modes of MTS, and how thermostats can be used for mitigation. Here we use the implementation in i-PI (Kapil et al., 2016; Litman et al., 2024), and use the non-conservative forces as the “fast” forces, $\mathbf{f}_{\text{fast}}=\mathbf{f}_{\text{NC}}$, and the difference between conservative and non-conservative forces as a slow correction, $\mathbf{f}_{\text{slow}}=\mathbf{f}_{\text{C}}-\mathbf{f}_{\text{NC}}$.
@@
Figure 13: A comparison of kinetic temperature profiles for NVE MD trajectories performed with the PET-M model and different orders of MTS integration. The step numbers refer to the “fast” evaluation.
@@
Appendix I Software
@@
ORB, MACE-MP-0, Equiformer, SevenNet calculations were performed using the publicly available repositories, respectively https://github.com/orbital-materials/orb-models/ (orb-models 0.4.0), https://github.com/ACEsuit/mace (mace-torch 0.3.8), https://github.com/FAIR-Chem/fairchem (fairchem-core 1.3.0), and https://github.com/MDIL-SNU/SevenNet (sevenn 0.10.1).

### 27 May 2025 v3

@@ 4.5 Equilibrium properties in the NVT ensemble
Judging by the average temperature $\langle T\rangle$ (Table 2), it is relatively easy to control the non-conservative behavior using a white-noise (WN) Langevin thermostat. However, strong couplings $\tau$ (the time scale over which the thermostat interferes with atomic motion) are needed, as even at 100 fs the average temperature is significantly off the target value. We discuss how these deviations in the equilibrium temperature affect structural properties of water in Section G.5. The upshot is that for accurate models and strong thermostatting, the effects are small but noticeable. Furthermore, the strong Langevin thermostatting reduces the sampling efficiency, and so even with a respectable 1 ns trajectory, many simple structural averages are not fully converged.
@@ 4.6 Sampling efficiency and time-dependent properties
Aggressive Langevin dynamics is bound to dramatically change time-dependent properties, and in particular to reduce the diffusion coefficient – so that longer trajectories are needed to collect statistically independent atomic configurations. This slow-down is apparent when looking at the velocity correlation spectra (Figure 3). In the weak coupling regime (WN, $\tau=1000$ fs) there is a (small) increase in diffusion coefficient relative to the reference, because of the unphysically higher temperature, while the high-frequency peaks corresponding to stretching and bending are only weakly perturbed. Stronger couplings alter the dynamics dramatically, and reduce the diffusion coefficient (and hence the efficiency in sampling slow, collective motion) by a factor of about 1.5 ($\tau=100$ fs) and 5 ($\tau=10$ fs), negating the inference speed-up of a non-conservative model – while making it impossible to accurately evaluate any time-dependent property.
Figure 3: Velocity power spectrum $\hat{c}_{vv}(\omega)$, for different PET models and thermostat types: A conservative (C) and non-conservative (NC) model using white-noise Langevin (WN) and stochastic velocity rescaling (SVR) thermostats.
@@
Some of the more sophisticated thermostats used to stabilize other types of MD approximations – such as carefully tuned generalized Langevin equations (Ceriotti et al., 2009; Morrone et al., 2011) – can also be used to enforce more aggressive temperature control with reduced dynamical disruption in conservative molecular dynamics, but they still modify the natural dynamical properties significantly, and they can fail catastrophically when used with non-conservative forces, as shown in Section G.6).
@@ 4.7 Geometry optimization
MD performed at very low temperature can be regarded as a form of geometry optimization. Our observations from MD suggest that sufficiently accurate non-conservative models should be able to reach reasonable, low-energy structures. We restrict ourselves to optimization algorithms based only on gradients, to avoid the complication of using inconsistent energies, as discussed in Section 3.3, comparing the FIRE (Bitzek et al., 2006) algorithm – that is similar in spirit to zero-temperature MD – and LBFGS (Liu & Nocedal, 1989) – a quasi-Newton algorithm that uses an approximation of the Hessian to accelerate convergence. Comparing different models on the task of optimizing a water snapshot from a MD configuration (Figure 7) shows that inaccurate non-conservative models, such as SOAP-BPNN-NC, fail catastrophically at geometry optimization, while more accurate models, such as PET-NC and ORB, can reach a locally stable configuration, especially using FIRE. We note, however, that non-conservative models are less stable when used with a Hessian-based method, with large fluctuations in the residual force that make it hard to define a stopping criterion. On a practical level, non-conservative forces are bound to make geometry optimization more fragile, and to require careful choice of the minimization algorithm and its convergence parameters, as we observe in Appendix F when comparing different general-purpose models.
@@ 4.8 Non-conservative forces as accelerators
While we have observed that conservative MLIPs are better suited for practical simulations, we suggest that hybrid models, which additionally support direct, non-conservative, force predictions, can be used for faster inference and training. Such models can be obtained by training both force heads jointly, as demonstrated in the PET-M model, or, more efficiently, by first training a non-conservative model and then fine-tuning its energy head to yield conservative forces. As shown in Figure 4 and further discussed in Appendix H, conservative fine tuning leads to the accuracy and physical correctness of conservative models at highly reduced training time.
Figure 4: Training curves for PET-C and PET-NC models, and the conservative fine-tuning of a hybrid PET-M model initialized (epoch marked with arrow) from the potential energy head of the PET-NC model.
In simulations, one can then use the conservative forces of such a hybrid model for validation, error monitoring and correction, and the direct forces for faster inference. A good example is to use multiple time-stepping (MTS) techniques (Tuckerman et al., 1992) for molecular dynamics, where the non-conservative forces are used to integrate the equations of motion, and the conservative forces are applied every $M$ steps as a correction. This reduces the theoretical overhead of a conservative trajectory from a factor of $F\approx 2$ to one of 1+(F1)/M1+(F-1)/M. The results using this technique in Table 2, Figure 2 and in Appendix I are essentially indistinguishable from fully conservative ones, using $M=8$, which leads to a small, approximately 2020% slowdown compared to a direct-force, non-conservative trajectory. Appendix I contains further explanation of the MTS technique, as well as more detailed results for MTS simulations using models trained on the water and OC20 datasets. The technique can be also successfully used for constant-pressure simulations, as shown in Section G.3.
@@ 5 Discussion
Code and data required to reproduce the results in this work are available on Zenodo at https://zenodo.org/records/14778891. An example of how to perform multiple-time-step dynamics with conservative and direct forces can be found at https://atomistic-cookbook.org/examples/pet-mad-nc/pet-mad-nc.html. More details on the software used are available in Appendix J.
Acknowledgements
The authors would like to thank Federico Grasselli and Niklas Schmitz for stimulating discussion, and Rafael Gomez-Bombarelli for a coffee-break conversation which inspired us to look into this problem. ML and MC acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme Grant No. 101001890-FIAMMA. FB and MC acknowledge support from the NCCR MARVEL, funded by the Swiss National Science Foundation (SNSF, grant number 182892) and from the Swiss Platform for Advanced Scientific Computing (PASC).
@@ Appendix B Theoretical computational cost
Appendix C Range of back-propagated and direct force models
Figure 5: A schematic representation of the interactions in a (ML) interatomic potential. (a) Contributions to an $i$-centered prediction of energy or force from three or more neighbors, labeled $j$ and $k$. (b) Contributions from a $j$-centered energy prediction to the force on the $i$-th atom, computed by back-propagation.
The use of a cutoff to restrict the range of interactions is ubiquitous in the construction of physics-based potentials, and is also an integral part of descriptor-based ML potentials (Musil et al., 2021). It is often argued that models that incorporate correlations between at least two neighbors of each central atom achieve an effective interaction range of twice the cutoff distance (Artrith et al., 2011) (see Figure 5a). A similar effect also applies to message-passing architectures (Nigam et al., 2022).
One can see clearly that this extension of the interaction range beyond the cutoff (or the receptive radius of the GNN) does not apply to the case in which forces are predicted directly. Considering for simplicity the case of a three-body model in which the potential contribution from each atomic environment $i$ can be written as a sum of a function of its distances with two neighbors $j$ and $k$, and the distance between the neighbors $r_{jk}$
$V_{i}=\sum_{j,k}v(r_{ij},r_{ik},r_{jk}),$ (10)
one sees that the dependency on interatomic distances greater than the cutoff is due to the relative position of atoms other than the central atom. Thus, a direct model that predicts a similar 3-body force
$\mathbf{f}_{i}=\sum_{j,k}\mathbf{f}(r_{ij},r_{ik},r_{jk}),$ (11)
or any other functional form limited to the neighbors of the $i$-th atom, only contains information on atoms within the cutoff. The dependency of $\mathbf{f}_{i}$ on the coordinates of far-away atoms is a consequence of the fact that the total energy is built as a sum over multiple centers. It is only through the terms of the form $\partial V_{j}/\partial\mathbf{r}_{i}$ that occur naturally when evaluating forces through back-propagation, than the force depends on the position of the neighbor’s neighbors. More generally, in a message-passing implementation, backpropagation ensures that force predictions benefit from an effective receptive radius that is twice that of atom-centered energy predictions.
A further concern for direct force models is that – at least for low-body-order models – atom-centered descriptors can be shown to have low resolving power, with pairs of distinct atomic environments having precisely the same representation (Pozdnyakov et al., 2020). For all known degeneracies, descriptors centered on other atoms allows distinguishing the structure as a whole, and as a consequence the total energy and interatomic forces can be still differentiated. This would not be the case for direct force predictions, which would fail completely to differentiate degenerate pairs when using atom-centered low-body-order models, and would be more sensitive to numerical instabilities for higher-order models.
In practice, we find consistent evidence of the practical impact of these considerations. As an example, Table 3 reports the accuracy of PET-C and PET-NC models, using 2 and 3 message passing layers. It can be seen that the accuracy of the direct force model benefits much more from the increase in the receptive radius.
Model 2 message-passing layers 3 message-passing layers
PET-C 20.8 18.6
PET-NC 32.8 24.3
Table 3: Test-set force MAE, in units of meV/Å, for conservative and non-conservative force-only PET models using 2 and 3 message-passing layers, respectively. Upon increasing the receptive radius, the accuracy of the NC model improves by around 30%, while the accuracy of the C model only improves by around 10%.
Appendix D Models and architectures
@@
ORB-v2 Non-equivariant, non-conservative model, trained on the Alexandria and MPtrj datasets.
@@
D.1 Timings of general-purpose models
@@
D.2 Accuracy and timings of water models
@@
Appendix E Non-conservative work
@@
Appendix F Geometry optimization
F.1 Optimization trajectories
@@
We discuss first the results for the FIRE algorithm (left panel in Figure 7). Despite the very different level of accuracy, the conservative SOAP-BPNN model and PET converge to a similar configuration (with PET forces saturating at about $10^{-3}$ eV/Å, due to numerical precision). The non-conservative PET-NC model also converges to a similar structure, indicating that a sufficiently accurate non-conservative model can indeed be used with a gradient-based structural optimization algorithm. The non-conservative SOAP-BPNN model, however, displays a catastrophic mode of failure, with the force never decreasing below 0.1 eV/Å, and the trajectory drifting off in a different direction without ever reaching a stable state. ORB converges to a very different structure than the other models, which might be due to the different reference DFT functional used for training.
@@ F.2 Failure rates
Appendix G Additional molecular dynamics results
G.1 NVE temperature profiles
@@
We also perform NVE MD trajectories on a wider range of chemical systems using the general-purpose potentials MACE-MP, SevenNet, Equiformer and ORB. Fig. 9 shows that the non-conservative models ORB and Equiformer exhibit a strong kinetic energy gain during the simulation of graphene, amorphous carbon and FCC aluminium, with ORB exhibiting an explosive behavior for graphene (possibly because it is outside the training set). The conservative models exhibit stable temperature fluctuations, even for amorphous carbon where the high distortion of the structure leads to higher initial energy and broader kinetic energy fluctuations. The order of magnitude of the drift is similar for systems with different degrees of chemical heterogeneity.
G.2 NVT
@@
Table 8: Average kinetic temperatures (in K) for different models and thermostatting schemes. For SVR, we show also the separate kinetic temperature computed only for hydrogen and oxygen atoms. All thermostat time constants are reported in units of fs.
@@
G.3 Direct stresses: NPH and NPT
In order to test the effects of direct stress prediction, which can be employed in constant-pressure simulations, we run MD in the NPH and NPT ensembles. In this case, as a means to disentangle the effects of non-conservative stresses from those of non-conservative forces, we run the “non-conservative” (NC) models using non-conservative stresses and conservative forces. For these simulations, we employ the PET-MAD general-purpose potential (Mazitov et al., 2025), whose version 1.1 is trained to provide both conservative and non-conservative stresses.
Figure 10: Length of the cubic cell edge, as a function of time, during NPH and NPT simulations of liquid water using the PET-MAD model, in conservative mode, non-conservative mode, and using multiple time stepping with $M=8$. The Bussi-Zykova-Parrinello barostat (Bussi et al., 2009) was used in all simulations.
Fig. 10 shows how the conservative runs produce reasonable cell volumes, both in the constant-enthalpy NPH ensemble and the isothermal-isobaric NPT ensemble. The non-conservative NPH run shows rapid and unphysical expansion of the cell, which is reflected in the rapid growth in the enthalpy of the simulation (which is instead conserved very well by the conservative NPH run). This “explosion” of the cell is prevented by the thermostat in the NPT ensemble, although the exact value of the cell parameter is inconsistent with the conservative NPT baseline. Finally, we show a MTS (multiple time stepping, see App. I) run using $M=8$, which correctly reproduces the cell volume of the conservative NPT simulation while only evaluating the more expensive conservative stresses every 8 time steps. This simple example illustrates how all the considerations that were made for non-conservative forces in the context of NVE/NVT simulations are also applicable to NPH/NPT simulations when using direct stress predictions.
G.4 Energy conservation by velocity rescaling
The easiest way to stabilize NVE molecular dynamics using direct forces is to rescale the velocities (or, equivalently, the momenta) at each step to enforce conservation of energy, as proposed in Bigi et al. (2025). While this leads to an improvement, in the sense that NVE trajectories are guaranteed to be stable, many fundamental issues remain. As an example, Fig. 11 compares traditional NVE MD with a conservative model to NVE MD with a non-conservative model, where the latter is stabilized by velocity rescaling. The conservative model shows kinetic temperatures close to the temperature at which the starting structure was equilibrated, both for the overall simulation and for individual atomic types. In contrast, the non-conservative run departs from the equilibration temperature and, more importantly, it shows much different average kinetic energies for atoms belonging to different chemical species. This behavior is in violation of the principle of equipartition of energy, and it is a manifestation of the fact that the target ensemble (NVE, in this case) is not being sampled correctly. This is very similar to what we observe with global thermostats, and is an analogous manifestation of the fact that enforcing the correct values of global quantities does not guarantee that all the parts of a simulation will equilibrate correctly, in the absence of a well-defined Hamiltonian.
Figure 11: Total and atom-type-resolved kinetic energy profiles for NVE simulations of liquid water using the conservative and non-conservative modes of the PET-MAD (Mazitov et al., 2025) potential. In the non-conservative case, energy conservation is enforced by velocity rescaling.
G.5 Structural correlations
@@
G.6 Generalized Langevin equations
The generalized Langevin equation (GLE) thermostat (Ceriotti et al., 2010) is a local stochastic thermostat that extends traditional Langevin dynamics by introducing a history-dependent friction term. The possibility of tuning the memory kernel of the friction and the noise allows tuning the effect of the thermostat, by effectively coupling different modes of the physical system to different friction timescales. Thanks to this construction, near-optimal damping can be achieved for a large range of characteristic frequencies, as opposed to the traditional Langevin thermostat, which only achieves optimal damping (and therefore sampling) for a narrow frequency range.
Among the many different applications of GLE thermostats, that of Ref. 50 appears particularly suited to contrast the temperature drift observed with the direct prediction of forces: by applying a high-pass filter to the noise, it allows to thermostat aggressively the fast degrees of freedom of the system, without disrupting the long-time dynamics and the sampling of the diffusive degrees of freedom. We use the same parameters that have been used successfully to control temperature drift caused by the integration errors for aqueous systems in Ref. 50, using $1/\gamma_{0}=83.33\,\textrm{ps}$, $1/\gamma_{\infty}=10\,\textrm{fs}$, and $\omega_{F}=300\,\textrm{ps}^{-1}$, where all GLE parameters refer to their definition in Morrone et al. (2011).
We probe the behavior of conservative and non-conservative models under GLE-thermostatted dynamics. Table 9 shows that this GLE, that could successfully control integration errors for an empirical water model, fails catastropically when used with non-conservative models. This can be explained by the observation that the near-diffusive modes, which are those that show the most non-conservative behavior according to the analysis in Figure 1, are almost not damped at all by this thermostat. Even though one could in principle design a low-pass filter, or use more aggressive parameters for this high-pass thermostat, the analysis performed in Ref. 50 indicates that such GLE would slow down long-time sampling, exactly as it is the case for a traditional Langevin thermostat. By reducing the sampling efficiency, this effect negates the advantages of direct force prediction.
Table 9: Average kinetic temperatures (in K) for different models and atomic species, for the molecular dynamics of water using a GLE thermostat.
Thermostat type GLE GLE GLE
Atoms all H O
[…]
SOAP-BPNN-NC (NC) $1.6\cdot 10^{8}$ $2.1\cdot 10^{8}$ $5.9\cdot 10^{7}$
[…]
Appendix H Conservative fine-tuning
@@
Appendix I Multiple time stepping
@@
where $\mathbf{q}$ and $\mathbf{p}$ indicate the vectors containing positions and momenta of all particles, and $\mathbf{m}$ the masses. Forces have to be recomputed if they are needed after a position update. In the limit of ideal time scale separation, the resulting dynamics (and the sampled ensemble for constant-temperature simulations) is consistent with $V_{\text{slow}}$, at a much-reduced cost: Letting $F$ denote the slowdown of “slow” over “fast” forces, and assuming that an evaluation of “slow” forces also yields the “fast” ones, MTS reduces the theoretical cost of well-behaved molecular dynamics from $F$ to 1+(F1)/M1+(F-1)/M. Choosing $M$ therefore requires a trade-off between accuracy and speed. Morrone et al. (2011) discusses the failure modes of MTS at high $M$, and shows how to use thermostats for mitigation.
Here we use the implementation in i-PI (Kapil et al., 2016; Litman et al., 2024), and use the non-conservative forces as the “fast” forces, $\mathbf{f}_{\text{fast}}=\mathbf{f}_{\text{NC}}$, and the difference between conservative and non-conservative forces as a slow correction, $\mathbf{f}_{\text{slow}}=\mathbf{f}_{\text{C}}-\mathbf{f}_{\text{NC}}$. Our current implementation does not reuse computation between the conservative and non-conservative forces, and so the practical overhead is given by $1+F/M$, with an empirical factor of $F\approx 2$ (see Appendix B). As seen in Table 10, empirical timings match the theoretical ratios rather closely.
Type Relative speed (O on Sc-Co-Al alloy) Relative speed (water) Relative speed (theoretical)
NC $1.0$ $1.0$ -
C $1.93$ $1.92$ $F=1.925$
MTS, $M=4$ $1.50$ $1.45$ $1.48$
MTS, $M=8$ $1.19$ $1.18$ $1.24$
MTS, $M=16$ $1.05$ $1.04$ $1.12$
Table 10: Empirically observed slowdown of conservative MD simulations for multiple time-stepping with different strides $M$, as well as a simulation where the conservative forces are evaluated at every step, compared with the theoretical cost $1+F/M$ with $F=1.925$. Timings are given for selected systems from the experiments in Sections 4.4 and G, using runs in the NVE ensemble.
@@
Figure 16: A comparison of kinetic temperature profiles for NVE MD trajectories of bulk water performed with the PET-M model and different orders of MTS integration. The step numbers refer to the “fast” evaluation.
@@
We also investigate the behavior of MTS for the more generic PET potential trained on the OC20 dataset in Appendix H. We use the model after conservative fine-tuning, which is able to predict both conservative and non-conservative forces with a good accuracy. The simulations were run on a system of a single oxygen atom adsorbed on an alloy surface, and their kinetic energy profiles are shown in Figure 18. Also in this case, well-behaved molecular dynamics can be performed up to a high multiple time-step factor $M=8$, a value that retains nearly all the speed-up afforded by non-conservative forces.
Figure 18: A comparison of kinetic temperature profiles for NVE MD trajectories of an oxygen atom adsorbed on an alloy surface, performed with the various heads of a model which was produced by “conservative fine-tuning” on a subset of the OC20 dataset. The step numbers refer to the “fast” evaluation.
Appendix J Software
@@
ORB, MACE-MP-0, Equiformer, SevenNet calculations were performed using the publicly available repositories, respectively https://github.com/orbital-materials/orb-models/ (orb-models 0.4.0), https://github.com/ACEsuit/mace (mace-torch 0.3.8), https://github.com/FAIR-Chem/fairchem (fairchem-core 1.3.0), and https://github.com/MDIL-SNU/SevenNet (sevenn $0.1$).